The pressure inside a fire hose is 1.6 x 10^5 Pa and the water inside moves at 0.68 m/s. At the nozzle, the pressure decreases to atmospheric pressure. Using Bernoulli's Principle, determine the velocity of the water exiting the nozzle if there is no change in height

1
Expert's answer
2022-01-30T14:18:38-0500

"p_1+\frac{\rho v_1^2}2=p_2+\frac{\rho v_2^2}2,"

"v_2=\sqrt{\frac 2 \rho(p_1-p_2)+v_1^2}=10.8~\frac ms."